# Lagrangian extensions and left symmetric structures on the four-dimensional real Lie superalgebras

### v1

@@ 6. The list of four-dimensional real Lie superalgebras
$D^{16}$ $\begin{array}[]{lcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None --
Table 4. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=1|3$
The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $p(\omega)$
$(A_{3,1}+A)$ $\begin{array}[]{lcrlcl}[e_{2},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{1}\end{array}$ None --
$\begin{array}[]{l}(D^{3}_{p,-1/2})\\[2.84526pt] p\not=0\end{array}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] \end{array}$ None --
$(D^{3}_{-1/2,-1/2})$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-\frac{1}{2}e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\\[2.84526pt] \end{array}$ $\begin{array}[]{l}\frac{1}{2}\lambda e_{1}^{*}\wedge e_{2}^{*}+\mu e_{1}^{*}\wedge e_{3}^{*}\\[2.84526pt] +\nu e_{1}^{*}\wedge e_{4}^{*}+\gamma e_{3}^{*}\wedge e_{4}^{*}\\[2.84526pt] +\frac{1}{2}\lambda e_{4}^{*}\wedge e_{4}^{*},\text{ where }\gamma\lambda\not=0\end{array}$ NH
$(D^{2}_{-1/2})^{1}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
$(D^{2}_{-1/2})^{2}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None --
Table 5. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=3|1$
The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $p(\omega)$
$(A_{1,1}+3A)^{1}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&e_{1}\end{array}$ None --
$(A_{1,1}+3A)^{2}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&-e_{1}\end{array}$ None --
Table 6. Trivial algebras (i.e. $[\mathfrak{g}_{\bar{1}},\mathfrak{h}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=1|3$
@@
Acknowledgment. We would like to acknowledge Chengming Bai for pointing out that Novikov algebras have two superizations. Also, we would like to thank Saïd Benayadi, Isabel Cunha, and Quentin Ehret for reading the manuscript and making pertinent comments.

### v2

@@ 6. The list of four-dimensional real Lie superalgebras
$D^{16}$ $\begin{array}[]{lcl}[e_{1},e_{2}]&=&e_{2},\\[2.84526pt] [e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&e_{3}+e_{4}\end{array}$ None -
Table 9. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]=\{0\}$) with $\mathrm{sdim}=1|3$
The LSA Relations in the basis: $e_{1},e_{2},e_{3}\,|\,e_{4}$ Symplectic structure $|\omega|$
$(A_{3,1}+A)$ $\begin{array}[]{lcrlcl}[e_{2},e_{3}]&=&e_{1},&[e_{4},e_{4}]&=&e_{1}\end{array}$ None -
[…]
$(D^{2}_{-1/2})^{1}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
$(D^{2}_{-1/2})^{2}$ $\begin{array}[]{lcrlcl}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{2}+e_{3},\\[2.84526pt] [e_{1},e_{4}]&=&\frac{1}{2}e_{4},&[e_{4},e_{4}]&=&e_{2}\end{array}$ None -
Table 10. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{g}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=3|1$
The LSA Relations in the basis: $e_{1}\,|\,e_{2},e_{3},e_{4}$ Symplectic structure $|\omega|$
$(A_{1,1}+3A)^{1}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&e_{1}\end{array}$ None -
$(A_{1,1}+3A)^{2}$ $\begin{array}[]{lcrlcl}[e_{2},e_{2}]&=&e_{1},&[e_{3},e_{3}]&=&e_{1},\\[2.84526pt] [e_{4},e_{4}]&=&-e_{1}\end{array}$ None -
Table 11. Trivial algebras (i.e. $\mathfrak{g}_{\bar{1}},\mathfrak{h}_{\bar{1}}]\not=\{0\}$) with $\mathrm{sdim}=1|3$
7. Appendix: The decomposable case
We are only considering Lie superalgebras that are not Lie algebras. We investigate symplectic structures on the decomposable 4-dimensional real Lie superalgebras below. The notation we use is that of Backhouse [B]. Over $A\oplus B$, the notation $\omega:=\omega_{A}\oplus\omega_{B}$ means that $\omega(a_{1}+b_{1},a_{2}+b_{2})=\omega_{A}(a_{1},a_{2})+\omega_{B}(b_{1},b_{2})$ for all $a_{1},a_{2}\in A$ and $b_{1},b_{2}\in B$.
7.1. Abelian LSA
  1. (1)
    NH: $\mathbb{R}^{1|3},$ $\mathbb{R}^{3|1},$
  2. (2)
    Symplectic: $(\mathbb{R}^{2|2},\omega_{\mathbb{R}^{1|1}}\oplus\omega_{\mathbb{R}^{1|1}})$, $(\mathbb{R}^{2|2},\omega_{\mathbb{R}^{2|0}}\oplus\omega_{\mathbb{R}^{0|2}})$, $(\mathbb{R}^{0|4},\omega_{\mathbb{R}^{0|2}}\oplus\omega_{\mathbb{R}^{0|2}}).$
7.2. LSA of the form $L\oplus M$, where $L,M$ are 2-dimensional
We refer to Table 12 for the description of the LSA below:
  1. (1)
    NH: $B\oplus\mathbb{R}^{2|0},$ $B\oplus\mathbb{R}^{0|2},$ $B\oplus\mathrm{Al},$ $\mathrm{Al}\oplus\mathbb{R}^{1|1},$
  2. (2)
    None: $(A_{1,2}+A)\oplus\mathbb{R}^{2|0}$, $(A_{1,2}+A)\oplus\mathbb{R}^{0|2},$ $(A_{1,2}+A)\oplus\mathbb{R}^{1|1}$, $(A_{1,2}+A)\oplus\mathrm{Al}$, $(A_{1,2}+A)\oplus B$, $(A_{1,2}+A)\oplus(A_{1,2}+A)$,
  3. (3)
    Symplectic: $(B\oplus B,\omega_{B}\oplus\omega_{B})$, $(B\oplus\mathbb{R}^{1|1},\omega_{B}\oplus\omega_{\mathbb{R}^{1|1}})$, $(\mathrm{Al}\oplus\mathbb{R}^{0|2},\omega_{\mathrm{Al}}\oplus\omega_{\mathbb{R}^{0|2}})$.
The LA Relations in the basis: $e_{1},e_{2}$ Symplectic structure $|\omega|$
$\mathrm{Al}$ $\begin{array}[]{lcr}[e_{1},e_{2}]&=&e_{1}\end{array}$ $\omega_{\mathrm{Al}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{0}$
$\mathbb{R}^{2|0}$ $\omega_{\mathbb{R}^{2|0}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{0}$
The LSA Relations in the basis: $e_{1}\,|\,e_{2}$ Symplectic structure $|\omega|$
$B$ $\begin{array}[]{lcr}[e_{1},e_{2}]&=&e_{2}\end{array}$ $\omega_{B}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{1}$
$(A_{1,2}+A)$ $\begin{array}[]{lcr}[e_{2},e_{2}]&=&e_{1}\end{array}$ None -
$\mathbb{R}^{1|1}$ $\omega_{\mathbb{R}^{1|1}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{1}$
The LSA Relations in the basis: $|\,e_{1},e_{2}$ Symplectic structure $|\omega|$
$\mathbb{R}^{0|2}$ $\omega_{\mathbb{R}^{0|2}}:=e_{1}^{*}\wedge e_{2}^{*}$ $\overline{0}$
Table 12. Quasi-Frobenuis structure
7.3. LSA of the form $L\oplus\mathbb{R}^{0|1}$ or $L\oplus\mathbb{R}^{1|0}$, where $L$ is 3-dimensional
Below, $X$ represents the generator of either $\mathbb{R}^{1|0}$ or $\mathbb{R}^{0|1}$. The parity of $X$ should be understood from the context.
The LSA Relations in the basis: $e_{1},e_{2}\,|\,e_{3}$ and $X$ Symplectic structure $|\omega|$
$C^{1}_{p}\oplus\mathbb{R}^{1|0},p\not=0$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\end{array}$ None -
$C^{1}_{p}\oplus\mathbb{R}^{0|1},p\not=0$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&pe_{3}\end{array}$ None -
$C^{1}_{-1}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3}\end{array}$ NH -
$C^{1}_{-1}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&-e_{3}\end{array}$ $e_{1}^{*}\wedge X^{*}+e_{2}^{*}\wedge e_{3}^{*}$ $\bar{1}$
$C^{1}_{1/2}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3},\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}\end{array}$ None -
$C^{1}_{1/2}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{2}]&=&e_{2},&[e_{1},e_{3}]&=&\frac{1}{2}e_{3},\\[2.84526pt] [e_{3},e_{3}]&=&e_{2}\end{array}$ $e_{3}^{*}\wedge e_{3}^{*}+X^{*}\wedge X^{*}$ $\bar{0}$
The LSA Relations in the basis: $e_{1}\,|\,e_{3},e_{4}$ and $X$ Symplectic structure $|\omega|$
$C^{2}_{p}\oplus\mathbb{R}^{1|0},\;0<|p|\leq 1$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&pe_{4}\end{array}$ None -
$C^{2}_{p}\oplus\mathbb{R}^{0|1},\;0<|p|\leq 1$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&pe_{4}\end{array}$ None -
$C^{2}_{-1}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&-e_{4}\end{array}$ $e_{1}^{*}\wedge e_{4}^{*}+X^{*}\wedge e_{4}^{*}$ $\bar{1}$
$C^{2}_{-1}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]&=&-e_{4}\end{array}$ NH -
$C^{3}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcr}[e_{1},e_{4}]&=&e_{3}\end{array}$ $e_{1}^{*}\wedge e_{3}^{*}+X^{*}\wedge e_{4}^{*}$ $\bar{1}$
$C^{3}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcr}[e_{1},e_{4}]&=&e_{3}\end{array}$ NH -
$C^{4}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]=e_{3}+e_{4}\end{array}$ None -
$C^{4}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&e_{3},&[e_{1},e_{4}]=e_{3}+e_{4}\end{array}$ None -
$C^{5}_{p}\oplus\mathbb{R}^{1|0},\;p\geq 0$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&pe_{3}-e_{4},&[e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None -
$C^{5}_{p}\oplus\mathbb{R}^{0|1},\;p\geq 0$ $\begin{array}[]{lcrlcr}[e_{1},e_{3}]&=&pe_{3}-e_{4},&[e_{1},e_{4}]&=&e_{3}+pe_{4}\end{array}$ None -
$(A_{1,1}+2A)^{1}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=e_{1}\end{array}$ None -
$(A_{1,1}+2A)^{1}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=e_{1}\end{array}$ None -
$(A_{1,1}+2A)^{2}\oplus\mathbb{R}^{1|0}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=-e_{1}\end{array}$ None -
$(A_{1,1}+2A)^{2}\oplus\mathbb{R}^{0|1}$ $\begin{array}[]{lcrlcr}[e_{3},e_{3}]&=&e_{1},&[e_{4},e_{4}]=-e_{1}\end{array}$ None -
Table 13. Quasi-Frobenuis structure
@@ 7.2. LSA of the form $L\oplus M$, where $L,M$ are 2-dimensional
Acknowledgment. We would like to acknowledge Chengming Bai for pointing out that Novikov algebras admit two superizations. Also, we would like to thank Saïd Benayadi, Isabel Cunha, Quentin Ehret, and the referee for reading the manuscript and making pertinent comments.
Declarations:
Funding Statement: SB’s research was supported by GRANT: AD-065.
Competing Interest Statement: the authors declare no competing interest.
Authors’ Contribution Statement: the authors contributed equally to this work.